The base of a triangle is 2 in.
longer than the height. The area is 12in^2 . Find the
height and the base.
How would I write this equation?

Answer by nerdybill(7384)

The base of a triangle is 2 in.
longer than the height. The area is 12in^2 . Find the
height and the base.
.
area for any triangle is:
(1/2)bh
where
b is base
h is height
.
Let h = height
then because "The base of a triangle is 2 in" we have
h+2 = base
.
The formula is:
(1/2)(h+2)h = 12
.
Solving, we start by multiplying both sides by 2 to get:
(h+2)h = 24
h^2 + 2h = 24
h^2 + 2h - 24 = 0
(h+6)(h-4) = 0
.
h = {-6, 4}
we can toss out the negative value leaving:
h = 4 inches (height)
.
Base:
h+2 = 4+2 = 6 inches